OpenStudy (anonymous):

Solve for x. 6x - 15 = 9x - 6 x = -7 x = -3 x = 1 ---- 3 x = 7 ------- 3

OpenStudy (misty1212):

\[6x-15=9x-6\]?

OpenStudy (misty1212):

a) subtract \(6x\) from both sides to get rid of them on the left b) add \(6\) to both sides and then c) divide both sides by \(3\)

OpenStudy (anonymous):

subtract 9x from both sides

OpenStudy (anonymous):

add 15 to both sides

OpenStudy (anonymous):

and divide both sides by -3

OpenStudy (anonymous):

or do what misty1212 said

OpenStudy (anonymous):

6x−15=9x−6 -9x -9x

OpenStudy (anonymous):

−3x−15=−6

OpenStudy (anonymous):

−3x−15=−6 + 15 +15

OpenStudy (anonymous):

−3x=9

OpenStudy (anonymous):

divide both sides by -3

OpenStudy (anonymous):

x=−3

OpenStudy (anonymous):

or do

OpenStudy (anonymous):

6x−15=9x−6 -6x -6x

OpenStudy (anonymous):

-15=3x-6

OpenStudy (anonymous):

-15=3x-6 +6 +6

OpenStudy (anonymous):

-9=3x

OpenStudy (anonymous):

divide by 3 on both sides....

OpenStudy (anonymous):

-3=x

OpenStudy (anonymous):

same answer